# MATHGLANCE: Multimodal Large Language Models Do Not Know Where to Look in Mathematical Diagrams

### v1

@@ Appendix C Problem Templates
We design three types of questions: multiple-choice, true/false, and open-ended. Specifclly, for multiple-choice questions, we incorporate plausible distractors to challenge the model’s geometric perception. These distractors are selected from geometric candidate pools, varying from visually similar to clearly distinguishable options, ensuring a balanced evaluation of fine-grained visual discrimination. By leveraging this template engine, MATHGLANCE ensures consistent question formulation, controlled variation in difficulty, and robust coverage of geometric primitives and their relationships, forming a reliable testbed for evaluating diagram perception in MLLMs.
@@
Appendix D Reformat Dataset for Solid Geometry & Graphs
@@
Appendix E Example Illustration of GeoPeP
Figs. 12-14 demonstrate how GeoPeP delivers explicit geometric information—covering object count, shape classification, fine-grained bounding box coordinates, and inter-object relationships—presented in both caption-style and instruction-following conversational formats.
Appendix F Visual Distractors
@@
By gradually increasing the Gaussian noise level, we can systematically evaluate its impact on the ability of MLLM to recognize mathematical structures. As shown in Fig. 11, the image transitions from a noise level of 0.1 to 0.8, progressively blurring the geometric features. We observe that as the noise intensity increases, the difficulty of recognizing mathematical structures also rises, posing greater challenges to accurate recognition and reasoning. When the noise level reaches $\sim\mathcal{N}(\mathbf{0},\mathbf{0.8})$, the mathematical structures become nearly indistinguishable to the human eye, resulting in a significant performance drop on MATHGLANCE, as illustrated in Fig. 10.
Appendix G Case Study
@@
Model Responses. In Fig. 15 and Fig. 16, we present a comparative analysis of the performance of different Multimodal Large Language Models (MLLMs), including SVE-Math-DeepSeek [55] (baseline), SVE-Math-DeepSeek+, and InterVL2.5 [10]. Our evaluation reveals that SVE-Math-DeepSeek+ achieves a significant improvement in response accuracy compared to its baseline, SVE-Math-DeepSeek. This improvement suggests that our carefully curated training data effectively enhances the model’s mathematical perception, enabling it to better comprehend problem structures, reason through mathematical concepts, and generate more precise answers. Furthermore, we observe that even InterVL2.5, a strong competitor in the field, produces incorrect responses in certain cases. Upon closer examination, these errors often stem from inaccurate mathematical perception rather than purely computational mistakes. This observation reinforces the notion that an MLLM’s ability to accurately perceive and interpret mathematical content is a critical factor in achieving high performance. Overall, our findings highlight the fundamental role of precise perception in mathematical reasoning, akin to its importance in vision and language understanding. Just as accurate perception is essential for tasks such as image recognition and natural language processing, it is equally vital for MATH-related problem-solving. Our study underscores that enhancing a model’s mathematical perception can lead to substantial gains in accuracy and reliability.
@@
Error Examples. In this section, we provide more detailed error examples of GPT-4o, Qwen2.5-VL-7B, and our 7B model. We categorize the errors into two types: Chain-of-Thought (CoT) Errors and Recognition Errors. CoT errors occur when the model engages in step-by-step reasoning for perception questions but ultimately provides incorrect answers. Recognition errors, on the other hand, arise when the model attempts direct answering without reasoning yet fails to produce the correct result. Representative examples for each error type are illustrated in Figs. 17-26.
Figure 12: Sampels of caption-style and instruction-following GeoPeP.
Figure 13: Sampels of caption-style and instruction-following GeoPeP.
Figure 14: Sampels of caption-style and instruction-following GeoPeP.
@@
Table 7: Examples of problem templates used by MATHGLANCE on different source data acorss different tasks.
@@
There {be} {num} {shape} in the picture, right? Choices: A:Yes B:No.
{be} there {num} {shape} in the picture? Choices: A:Yes B:No.
You can see {num} objects in the picture, can’t you? Choices: A:Yes B:No.
There should be {num} shapes in the picture, correct? Choices: A:Yes B:No.
@@
Can the relationship {preposition} {shape} in the image be described as ”{relation}”? Choices: A:Yes B:No.
Does the image show the relationship {preposition} {shape} as ”{relation}”? Choices: A:Yes B:No.
@@
There {be} {num} {shape} in the picture, right? Choices: A:Yes B:No.
{be} there {num} {shape} in the picture? Choices: A:Yes B:No.
There are {num} objects in the picture, right? Choices: A:Yes B:No.
Are there {num} shapes in the picture? Choices: A:Yes B:No.
@@
{be} there {num} {shape} in the picture? Choices: A:Yes B:No.
The picture contains {num} {shape}, each represented by a different color, doesn’t it? Choices: A:Yes B:No.
You can see {num} {shape} with different colors in the picture, can’t you? Choices: A:Yes B:No.
@@
Figure 15: Response comparisons between SVE-Math-DeepSeek $\!\!{}^{+}$, SVE-Math-DeepSeek, and InterVL2.5.
Figure 16: Response comparisons between SVE-Math-DeepSeek $\!\!{}^{+}$, SVE-Math-DeepSeek, and InterVL2.5.

### v2

@@ Appendix E Problem Templates
We design three types of questions: multiple-choice, true/false, and open-ended. Specifclly, for multiple-choice questions, we incorporate plausible distractors to challenge the model’s geometric perception. These distractors are selected from geometric candidate pools, varying from visually similar to clearly distinguishable options, ensuring a balanced evaluation of fine-grained visual discrimination. By leveraging this template engine, MATHEMETRIC ensures consistent question formulation, controlled variation in difficulty, and robust coverage of geometric primitives and their relationships, forming a reliable testbed for evaluating diagram perception in MLLMs.
@@
Appendix F Dataset Construction for Plane/Solid Geometry & Graphs
F.0.1 Synthetic Data Engine for Plane Geometry
Fig. 4 (main paper) describes the entire generation process. Inspired by AlphaGeometry [26], we use geometric clauses as fundamental units to construct complex plan geometric figures. A geometric clause is a formalized description of basic geometric objects and mathematical relationships, along with their properties or attibutes, a.k.a. prerequisite points. We first construct two geometry substrate pools, one containing 16 different geometric shapes (i.e., isosceles triangle, square, rectangle, parallelogram, isosceles trapezoid, pentagon, circle $\dots$ ellipse and segment), and the other defining 10 mathematical relationships (i.e., on, intersection, parallel, perpendicular, tangent, $\dots$ and reflection). We then randomly sample one or more substrates from these pools and pass them through a verifier, which ensures that logically paired shapes and relationships are preserved in the construction of valid geometric images. The verifier makes decisions based on either manually designed rules, fundamental mathematical knowledge, or prerequisite points. For example, parallel lines cannot intersect, and without enough prerequisite points, an angle trisection clause is invalid. In general, the chosen relationships are enforced by introducing additional shapes into the figure, ensuring that the specified relationships are accurately maintained and geometrically consistent throughout the construction process. We finally save the outputs as structured JSON annotations for image rendering using the Matplotlib package [27] and generating question-answer pairs via template-based pipelines.
Image Generation. The JSON annotations define foreground and background styles, i.e., colors sampled from a monochromatic palette, line width, and font size, as well as object shape information (class names), attributes (vertices labeled with random letters), bounding box locations, and mathematical relationships. Notably, spatial relationships are generated based on the bounding box locations of two objects (e.g., top-left, top-right, bottom-left, bottom-right). Similar to the image rendering process in [26], we translate geometric clauses into visual representations using Python code. This process first determines coordinates of points defined in each clause for basic shape visualization, then generates new primitives based on specified relationships. To increase task difficulty, we apply image augmentations, e.g., Gaussian noise, irregular scribbles, and wedge-shaped symbols for congruent angles or auxiliary lines (see Figs. 10 and 12).
Question-Answer Generation. Based on structured annotations, we employ a template-based pipeline to generate multiple-choice and true-false questions across four perception tasks. For example, in constructing Q&A pairs for shape classification, we first randomly select an object and its associated attributes from the given figure. We then formulate a question by filling placeholders in our carefully designed templates. For true-false questions, the ground truth answer is directly derived from the JSON annotations. For multiple-choice questions, we generate plausible distractors to challenge the model’s geometric perception, combining them with the correct answer into a set of four choices. Distractors are selected from geometric candidate pools, ranging from visually similar (e.g., equilateral vs. isosceles triangles) to visually dissimilar pairs (e.g., equilateral triangle vs. circle). Further details on the Q&A generation templates are provided in Tab. 13.
Copyright and Consent: For GEOMETRIC, we will release the entire package including all data and code used to generate samples under CC BY 4.0 on acceptance.
F.0.2 Reformat Dataset for Solid Geometry & Graphs
Copyright and Consent: CLEVR [24] is under CC BY 4.0 and FigureQA [25] is under CC0 1.0. All files that we re-publish (JSON reformats and look-up tables) will carry the same licenses as their sources.
@@ F.0.2 Reformat Dataset for Solid Geometry & Graphs
Appendix G Example Illustration of GEOMETRIC
Figs. 13-15 demonstrate how GEOMETRIC delivers explicit geometric information—covering object count, shape classification, fine-grained bounding box coordinates, and inter-object relationships—presented in both caption-style and instruction-following conversational formats.
Appendix H Visual Distractors
@@
By gradually increasing the Gaussian noise level, we can systematically evaluate its impact on the ability of MLLM to recognize mathematical structures. As shown in Fig. 12, the image transitions from a noise level of 0.1 to 0.8, progressively blurring the geometric features. We observe that as the noise intensity increases, the difficulty of recognizing mathematical structures also rises, posing greater challenges to accurate recognition and reasoning. When the noise level reaches $\sim\mathcal{N}(\mathbf{0},\mathbf{0.8})$, the mathematical structures become nearly indistinguishable to the human eye, resulting in a significant performance drop on MATHEMETRIC, as illustrated in Fig. 11.
Appendix I Case Study
@@
Model Responses. In Figs. 16-18, we present a comparative analysis of model responses across several Multimodal Large Language Models (MLLMs), including SVE-Math-DeepSeek [19], InternVL2.5 [35], Qwen2.5-VL [18], and GPT-4o, alongside our enhanced variants: SVE-Math-DeepSeek+ and Qwen2.5-VL+. Our evaluation shows that fine-tuning base models (SVE-Math-DeepSeek and Qwen2.5-VL) on GEOMETRIC significantly improves response accuracy over their respective baselines. This improvement suggests that our carefully curated training data effectively enhances the model’s mathematical perception, enabling it to better comprehend problem structures, reason through mathematical concepts, and generate more precise answers. For example, in Fig 16 (a), the base model fails to recognize the hypotenuse, while the GEOMETRIC model correctly identifies it and naturally applies the Pythagorean theorem to reach the correct answer. Similarly, in Fig 16 (b), the GEOMETRIC model correctly perceives the diameter, infers the presence of a right triangle, and again applies the appropriate theorem. In Fig 18 (a), both the base model and ours know the relevant theorem—the Inscribed Angle Theorem. However, the base model misperceives key visual cues (e.g., tangent lines), leading to hallucinated angle relations and an incorrect final answer. The GEOMETRIC model correctly perceives the tangency and the right-angle structure, enabling it to execute the theorem correctly and arrive at the correct answer. Furthermore, we observe that even GPT-4o, a strong competitor in the field, produces incorrect responses in certain cases. Upon closer examination, these errors often stem from inaccurate mathematical perception rather than purely computational mistakes. This observation reinforces the notion that an MLLM’s ability to accurately perceive and interpret mathematical content is a critical factor in achieving high performance. Overall, our findings highlight the fundamental role of precise perception in mathematical reasoning, akin to its importance in vision and language understanding. perception can lead to substantial gains in accuracy and reliability.
@@
Error Examples. In this section, we provide more detailed error examples of GPT-4o, Qwen2.5-VL-7B, and our 7B model. We categorize the errors into two types: Chain-of-Thought (CoT) Errors and Recognition Errors. CoT errors occur when the model engages in step-by-step reasoning for perception questions but ultimately provides incorrect answers. Recognition errors, on the other hand, arise when the model attempts direct answering without reasoning yet fails to produce the correct result. Representative examples for each error type are illustrated in Figs. 19-28.
Appendix J Limitations and Broader Impacts
Limitations. The proposed MATHEMETRIC focuses on evaluating MLLMs’ perceptual capabilities in structured, symbolic mathematical diagrams spanning plane geometry, solid geometry, and graphical representations. It enables fine-grained assessment through carefully designed perception tasks. Our findings clearly show that current models exhibit limited perceptual ability in diagram understanding across all three subjects, but our proposed training dataset, GEOMETRIC, remains limited to synthetic and geometry-based content. As such, it may not fully capture the diversity and ambiguity of real-world educational materials (e.g., textbook figures, handwritten notes, or scanned diagrams). In future work, we aim to extend our graph-based data construction to solid and graphical domains, which requires further investigation into their distinct structural representations. For instance, in graphical images such as line and bar charts, the underlying structure differs significantly from geometric diagrams. Nodes may correspond to data points, axis labels, or bars, while edges may capture trends, groupings, or relational mappings across axes. Developing meaningful node–edge representations in these contexts will be essential for enabling accurate visual grounding and interpretation. Moreover, we believe additional tasks, such as visual marker detection, are also critical for comprehensively evaluating MLLMs. For instance, recognizing special diagrammatic markers that represent relationships like parallelism or perpendicularity (e.g., double lines or small squares) is essential for deeper geometric understanding. Similarly, OCR capabilities are important, such as detecting angle measures or textual annotations embedded within diagrams. Expanding the benchmark to cover these aspects would further advance the assessment of MLLMs’ diagram perception abilities.
Furthermore, although we analyze how perception affects reasoning, our current setup does not explicitly model the interleaved interaction between visual grounding and logical reasoning—a critical direction for advancing multimodal mathematical reasoning.
Broader Impacts. This paper sheds light on a critical yet underexplored capability of MLLMs—mathematical diagram perception—and aims to push the field toward more interpretable and reliable multimodal reasoning. By isolating perception from reasoning, we reveal fundamental limitations in current models and provide a diagnostic tool for future model development. These insights could benefit applications in math education, AI-assisted learning, and cognitive science research. Caution is warranted, however, as over-reliance on synthetic datasets may lead to overfitting, and enhanced perceptual accuracy does not inherently ensure reliable reasoning. Furthermore, improved performance on benchmark tasks may not directly translate to robust generalization in real-world scenarios. We advocate for future research that bridges synthetic benchmarks with naturalistic data and advances human-aligned evaluation frameworks for visual reasoning systems.
Appendix K Use of Large Language Models (LLMs)
We used LLMs solely as an auxiliary tool for grammar correction and minor language polishing. They did not contribute to research ideation, experimental design, data analysis, or writing of technical content.
Figure 13: Sampels of caption-style and instruction-following GEOMETRIC.
Figure 14: Sampels of caption-style and instruction-following GEOMETRIC.
Figure 15: Sampels of caption-style and instruction-following GEOMETRIC.
@@
Table 13: Examples of problem templates used by MATHEMETRIC on different source data acorss different tasks.
@@
There {be} only {num} {shape} in the picture, right? Choices: A:Yes B:No.
{be} there only {num} {shape} in the picture? Choices: A:Yes B:No.
You can only see {num} objects in the picture, can’t you? Choices: A:Yes B:No.
There should be only {num} shapes in the picture, correct? Choices: A:Yes B:No.
@@
Can the relationship {preposition} {shape} in the image be described as "{relation}"? Choices: A:Yes B:No.
Does the image show the relationship {preposition} {shape} as "{relation}"? Choices: A:Yes B:No.
@@
There {be} only {num} {shape} in the picture, right? Choices: A:Yes B:No.
{be} there only {num} {shape} in the picture? Choices: A:Yes B:No.
There are only {num} objects in the picture, right? Choices: A:Yes B:No.
Are there only {num} shapes in the picture? Choices: A:Yes B:No.
@@
{be} there only {num} {shape} in the picture? Choices: A:Yes B:No.
The picture contains only {num} {shape}, each represented by a different color, doesn’t it? Choices: A:Yes B:No.
You can only see {num} {shape} with different colors in the picture, can’t you? Choices: A:Yes B:No.
@@
Figure 16: Response comparisons between SVE-Math-DeepSeek ${}^{+}$ (7B), SVE-Math-DeepSeek (7B), and InterVL2.5 (8B) in MathVerse.
Figure 17: Response comparisons between SVE-Math-DeepSeek ${}^{+}$ (7B), SVE-Math-DeepSeek (7B), and InterVL2.5 (8B) in MathVerse.
Figure 18: Response comparisons between Qwen2.5-VL-7B+ (Qwen2.5-VL-7B fine-tuned with GEOMETRIC using LoRA), GPT-4o, and Qwen-2.5-VL-7B in MathVerse.